OpenStudy (anonymous):

Find the value of tan (alpha–beta) if cos alpha= –3/5, sin beta= 5/13, 90° < alpha < 180°, and 90° < beta < 180°. ? HELP ANYONE!!!!

OpenStudy (tkhunny):

Do you have the formula for \(\tan(\alpha - \beta)\)? Note: 3, 4, and 5 constitute a Pythagorean Triple. \(3^{2} + 4^{2} = 5^{2}\) Note: 5, 12, and 13 constitute a Pythagorean Triple. \(5^{2} + 12^{2} = 13^{2}\)

OpenStudy (anonymous):

Look tan(a-b)=(tana-tanb)/1+tanatanb-----then tana=sina/cosb tanb=sinb/cosb After substitute the value of tana &tanb I left the rest for u!!!go Hurry up!!! a=alpha,b=beta